# Non-perturbative renormalization group for Higgs-like models in 4D

### v2

@@ VI.1 Massive Flows with fixed points
VI.2 Massless Flows between two non-trivial fixed points for imaginary coupling
@@
Thus at small coupling, the flows are no longer hyperbolas, but circles, and can thus can both originate and terminate along the line of fixed points. The trajectories based on the 3-loop $Q$ (67) are shown in Figure 11.
@@
$33<\Gamma_{\scriptscriptstyle{\rm UV}}<4~{}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}~{}0<g_{3}<4/$ 3 . (82)
@@
VI.3 Some special rational points in the 4D CFT’s
We have already seen that the point $g_{3}=4/3$ is special since it is at the boundary of the massless flows where $\Gamma_{\scriptscriptstyle{\rm IR}}=\infty$. One can surmise other potentially interesting special points as follows. When $\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D}$ where $\mathfrak{D}$ is an integer $\mathfrak{D}\geq 4$, then the IR theory has some likely interpretations. For instance, if $\mathfrak{D}=8$, this could represent massless flows that arrive to the IR via the operator $T_{\mu\nu}T^{\mu\nu}$, or $(T_{\mu}^{\mu})^{2}$ where $T_{\mu\nu}$ is the dimension 4 energy-momentum tensor. If this were the case, then this is analogous to $T\overline{T}$ perturbations of 2D CFT SmirnovZamo . In this case $\Gamma_{\scriptscriptstyle{\rm UV}}=16/5$ which is rational and corresponds to $g_{3}=1$, which is in the range (82). If one repeats this argument for the same beta function in 2 spacetime dimensions based on (100), one finds $\Gamma_{\scriptscriptstyle{\rm UV}}=4/3$. This turns out to be the scaling dimension of the perturbation for the minimal ${\cal N}=2$ supersymmetric model at Virasoro central charge $c=1$ BLsusy , which is a special point of the sine-Gordon model at coupling $b^{2}=4/3$ in (96). Thus the 2D analog of this flow is such that it preserves the central charge $c$, but the ${\cal N}=2$ supersymmetry is broken in the flow and the remaining massless fields are goldstinos. [^9] To our knowledge this observation is absent from the literature. This flow is similar to massless flows that arrive in the IR to the CFT of a Majorana fermion via the dimension 4 operator $T\overline{T}$ Alyosha , where the UV central charge is $7/10$. It was shown in AhnLeClair that the only UV completions of flows that end at the Majorana description of the Ising model have UV central charge $c=7/10$ and $c=3/2$, both of which have ${\cal N}=1$ supersymmetry. [^10] There are other flows that arrive to the Ising model via $T\overline{T}$ based on a different spectrum related to the $E_{8}$ Lie algebra AhnLeClair , indicating multi-critical behavior. In the notation in Tanaka , this is a flow between $c<1$ minimal models $M(11,12)\to M(3,4)$, with central charge $c_{\scriptscriptstyle{\rm UV}}=21/22$, $c_{\scriptscriptstyle{\rm IR}}=1/2$. This flow is not included in Tanaka since there the focus on the generic SU(2) quantum group modular fusion category of Verlinde lines, whereas our $E_{8}$ based flow should be understood from the $E_{8}$ quantum group fusion category.
@@ VI.3 Some special rational points in the 4D CFT’s
$\{\Gamma_{\scriptscriptstyle{\rm UV}},~{}\mathfrak{D}=4,5,6,7,8\}$ $=$ $\{4,\tfrac{7}{2},\tfrac{10}{3},\tfrac{13}{4},\tfrac{16}{5}\}$ (84)
$\{g_{3}^{\scriptscriptstyle{\rm UV}},~{}\mathfrak{D}=4,5,6,7,8\}$ $=$ $\{0,\tfrac{4}{7},\tfrac{4}{5},\tfrac{12}{13},1\}.$ (85)
@@ VI.4 Cyclic RG flows
Figure 12: Non-perturbative flows that are cyclic, based on the contour plot of the non-perturbative RG invariant $Q$ in equation (71) in the cyclic regime with $350.1<Q<$ 35. The points $g_{3}=\pm\infty$ are identified such that the topology of coupling constant space is that of a cylinder. Nowhere does the RG trajectory reach the fixed points along the $g_{3}$ axis.
Cyclic flows have previously been considered as rare, or just curiosities, because of various c-theorems and other considerationsctheorem ; CardyCtheorem ; Osborn ; JackOsborn ; Schwimmer ; Komargodski ; Polchinski . The latter studies often utilize quantum fields coupled to gravity since the focus is on the conformal, or trace anomaly of the stress-energy tensor. All of these works assume the theory is unitary. To address this further, first and foremost, our model certainly exists by construction, and as we showed, it has a regime with massless flows between conformal fixed points, which are more conventional flows. Although the CFT’s are non-unitary, these massless flows still fit in the paradigm that all flows begin and end at conformal fixed points. One should mention that in recent studies of massless flows between non-unitary minimal CFT’s in 2D, such flows are still consistent with a c-theorem Ravanini ; MussardoLG ; Tanaka ; KlebanovMM ; Negro , where here $c=c_{{\rm eff}}$, where $c_{{\rm eff}}$ is the effective central charge that appears in the thermodynamic Bethe ansatz. Returning to our model, if one is forced to understand the flows for all regimes of the couplings $g_{1},g_{3}$, then a cyclic regime where $Q$ is positive is an inescapable conclusion. At small coupling, this is just above the SU(2) symmetric diagonals. Secondly, for the 2D models with beta functions that are identical to those considered here and reviewed in the Appendix, an exact S-matrix was even proposed LeClairSierra , further justifying that these models are well-defined.
@@ VII Conclusions
We have computed the beta functions of the model in ALDolls to 3 loops, and showed that they do not spoil the 1-loop features found previously. For SU(2) broken to U(1) we re-summed an infinite number of contributions in order to obtain a non-perturbative beta function. This model has a line of fixed points, which are new non-unitary CFT’s in 4 spacetime dimensions. There exists RG flows between non-trivial fixed points and we computed the relation between the anomalous dimensions in the UV and IR, and furthermore identified some special points where these anomalous dimensions are rational. There also exists a regime where the flows do not begin nor end at a fixed point but in fact are cyclic, and we computed the period $\lambda$ of the flow in terms of the RG invariant $Q$. We argued that this circumvents the paradigm that all QFT’s begin or end at a fixed point as proposed in Schwimmer ; Polchinski , simply because the model is non-unitary.
@@
The analysis of the higher order RG flows in this paper points to the existence of some new non-unitary CFT’s in 4 spacetime dimensions. We computed some anomalous dimensions of relevant perturbing operators in the UV and based on massless flows found some special rational exponents corresponding to flows that arrive in the IR via operators of integer dimension $\mathfrak{D}>4$. If these calculations are correct, they are still far from a complete understanding of these potential CFT’s: what are the primary fields and their correlation functions? It is interesting to observe that rational exponents can exist in both 2 and 4 spacetime dimensions, whereas in 3 dimensions, non-trivial exponents obtained by means of the epsilon expansion are highly non-perturbative and irrational WilsonFisher . This is perhaps due to there being no obvious analog to the $J^{a}\widetilde{J}^{a}$ factorization of marginal operators in $3D$, which would require $J^{a}$ to have dimension $3/2$ such as $\phi^{3}$, such that the marginal interaction goes as $\phi^{6}$. In 2D the obvious generalization of this factorization is as a product of left-right moving chiral conserved currents, which have dimension 1.
Our models may have applications to statistical mechanics in 4 spatial dimensions where there are less restrictive positivity constraints. However we argued that the pseudo-hermitian structure of the hamiltonian, namely $H^{\dagger}={\cal K}H{\cal K}$ with ${\cal K}^{2}=1$, allows a consistent projection onto positive norm states, and can perhaps still describe consistent unitary quantum mechanics for QFT in $3+1$ spacetime dimensions. The modified inner product based on the operator ${\cal K}$ is essential to this construction.
@@
It would be clearly desirable to have a more fundamental understanding of the ${\cal K}$-insertions for our modified metric on the Hilbert space, since in this article it was just introduced in order to obtain the interesting OPE (18) since this structure leads to interesting beta-functions. Based on our results on massless flows above, by analogy with the classification of massless flows between non-unitary CFT;s in 2D Tanaka , we suggest that ${\cal K}$ can be interpreted as a topological defect, analogous to Verlinde lines in 2D CFT. In the 4D context, these should be codimension 2 defects, namely two dimensional surfaces. It is well accepted that such defects can modify the critical behavior in significant ways. The massless flows we found along the critical line $g_{1}=0$ should then be viewed as constrained by the U(1) fusion category. From this viewpoint, the line of fixed point we found are analogous to the radius of compactification of a free boson in 2D, which also has a duality. The flows in Tanaka are based on a study of the $SU(2)_{q}$ quantum group modular fusion category where $q$ is a root of unity, which is known to be a reduction of the U(1) category from various viewpoints. Very recent work of Witten is also suggestive of a such an understanding, where he argues that bras and kets can differ by an operator, which is analogous to our ${\cal K}$-inner product where kets have an operator ${\cal K}$ insertion, whereas bras do not.
@@
There is some evidence for an underlying ${\rm SL}(2,\mathbb{Z})$ structure in our model. Above we showed that the higher order beta functions have a $g\to 1/g$ duality, like an S-duality $\tau\to-1/\tau$. T-duality, $\tau\to\tau+1$ typically corresponds to a periodicity due to a $\vartheta$-angle, as in ${\cal N}=4$ supersymmetric Yang-Mills theory. The cyclicity of the RG required us to identify $g_{3}=\pm\infty$, endowing the 2-parameter coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder, which is a form of periodicity. This by itself does not establish an ${\rm SL}(2,\mathbb{Z})$ symmetry, since we have not precisely identified the elliptic modulus $\tau$ as a function of the 2 couplings. The role of ${\rm SL}(2,\mathbb{Z})$ is perhaps clearer for the fully anisotropic case with the beta functions (53) since integration of these 1-loop beta functions already leads to elliptic functions LeClairSierra . Furthermore the 2D models, with identical beta functions to our models, were proposed to have the S-matrix of Zamolodchikov Zelip built out of elliptic functions Elliptic . This S-matrix is characterized by a $\mathbb{Z}_{4}$ symmetry which is a subgroup of ${\rm SL}(2,\mathbb{Z})$, generated by the matrix $A=-i\sigma^{2}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Furthermore, it is known that this S-matrix transforms covariantly under ${\rm SL}(2,\mathbb{Z})$, meaning the S-matrix transforms with phase factors or gauge transformations that preserve the structure of the Yang-Baxter equation Baxter ; Belavin . In Elliptic , for the SU(2) broken to U(1) model, which is the cyclic sine-Gordon model reviewed in the Appendix, the latter can be obtained from certain limits of Zamolodchikov’s elliptic S-matrix, and this is perhaps a way of understanding what is left of the ${\rm SL}(2,\mathbb{Z})$ structure for our two parameter model.
VIII Acknowledgements
@@
We wish to thank Luis Alvarez-Gaumé, Denis Bernard, Igor Klebanov, Michael Peskin, Germán Sierra and Takahiro Tanaka for discussions.
IX Appendix: Comparison with 2D current-current perturbations
@@
It is instructive to compare the above 3-loop results to exactly solvable models in 2 spacetime dimensions which have identical 1-loop beta functions, since this can provide insights into higher orders in perturbation theory, and more importantly, it provides a guide as to what the physical implications could be.
@@
These flows only exist for 1¡$\Gamma_{\scriptscriptstyle{\rm UV}}<2$, where $\Gamma_{\scriptscriptstyle{\rm UV}}=1$ corresponds to the free-fermion point of the sine-Gordon model. This model is integrable and the relation (100) agrees exactly with the thermodynamic Bethe-ansatz analysis in Fendley2 , and this provides a non-trivial check of the validity of our beta functions to all orders.

### v3

@@ V.1 Massive Flows with fixed points
V.2 Massless Flows between two non-trivial fixed points for imaginary coupling
@@
Thus at small coupling, the flows are no longer hyperbolas, but circles, and can thus can both originate and terminate along the line of fixed points $g_{1}=0$. The trajectories based on the 3-loop $Q$ (50) are shown in Figure 11.
@@
$3<\Gamma_{\scriptscriptstyle{\rm UV}}<4~~~~~\Longrightarrow~~~~0<g_{3}<4/3.$ (65)
@@
V.3 Some special rational points in the 4D CFT’s
We have already seen that the point $g_{3}=4/3$ is special since it is at the boundary of the massless flows where $\Gamma_{\scriptscriptstyle{\rm IR}}=\infty$. One can surmise other potentially interesting special points as follows. When $\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D}$ where $\mathfrak{D}$ is an integer $\mathfrak{D}\geq 4$, then the IR theory has some likely interpretations. For instance, if $\mathfrak{D}=8$, this could represent massless flows that arrive to the IR via the operator $T_{\mu\nu}T^{\mu\nu}$, or $(T_{\mu}^{\mu})^{2}$ where $T_{\mu\nu}$ is the dimension 4 energy-momentum tensor. If this were the case, then this is analogous to $T\overline{T}$ perturbations of 2D CFT SmirnovZamo . For our model, $\Gamma_{\scriptscriptstyle{\rm UV}}=16/5$ when $\Gamma_{\scriptscriptstyle{\rm IR}}=8$, which is rational and corresponds to $g_{3}=1$, which is in the range (65). If one repeats this argument for the same beta function in 2 spacetime dimensions based on (83), one finds $\Gamma_{\scriptscriptstyle{\rm UV}}=4/3$. This turns out to be the scaling dimension of the perturbation for the minimal ${\cal N}=2$ supersymmetric model at Virasoro central charge $c=1$ BLsusy , which is a special point of the sine-Gordon model at coupling $b^{2}=4/3$ in (79). Thus the 2D analog of this flow is such that it preserves the central charge $c$, but the ${\cal N}=2$ supersymmetry is broken in the flow and the remaining massless fields are goldstinos. [^9] To our knowledge this observation is absent from the literature. This flow is similar to massless flows that arrive in the IR to the CFT of a Majorana fermion via the dimension 4 operator $T\overline{T}$ Alyosha , where the UV central charge is $7/10$. It was shown in UVIsing ; AhnLeClair that the only UV completions of flows that end at the Majorana description of the Ising model have UV central charge $c=7/10$ and $c=3/2$, both of which have ${\cal N}=1$ supersymmetry. [^10] There are other flows that arrive to the Ising model via $T\overline{T}$ based on a different spectrum related to the $E_{8}$ Lie algebra AhnLeClair , indicating multi-critical behavior. In the notation in Tanaka , this is a flow between $c<1$ minimal models $M(11,12)\to M(3,4)$, with central charge $c_{\scriptscriptstyle{\rm UV}}=21/22$, $c_{\scriptscriptstyle{\rm IR}}=1/2$. This flow is not included in Tanaka since there the focus on the generic SU(2) quantum group modular fusion category of Verlinde lines, whereas our $E_{8}$ based flow should be understood from the $E_{8}$ quantum group fusion category.
@@ V.3 Some special rational points in the 4D CFT’s
$\{\Gamma_{\scriptscriptstyle{\rm UV}}~{\rm for}~\mathfrak{D}=4,5,6,7,8\}$ $=$ $\{4,\tfrac{7}{2},\tfrac{10}{3},\tfrac{13}{4},\tfrac{16}{5}\}$ (67)
$\{g_{3}^{\scriptscriptstyle{\rm UV}}~{\rm for}~\mathfrak{D}=4,5,6,7,8\}$ $=$ $\{0,\tfrac{4}{7},\tfrac{4}{5},\tfrac{12}{13},1\}.$ (68)
@@ V.4 Cyclic RG flows
Figure 12: Non-perturbative flows that are cyclic, based on the contour plot of the non-perturbative RG invariant $Q$ in equation (54) in the cyclic regime with $0.1<Q<35$. The points $g_{3}=\pm\infty$ are identified such that the topology of coupling constant space is that of a cylinder. Nowhere does the RG trajectory reach the fixed points along the $g_{3}$ axis.
Cyclic flows have previously been considered as rare, or just curiosities, because of various c-theorems and other considerationsctheorem ; CardyCtheorem ; Osborn ; JackOsborn ; Schwimmer ; Komargodski ; Polchinski . The latter studies often utilize quantum fields coupled to gravity since the focus is on the conformal, or trace anomaly of the stress-energy tensor. All of these works assume the theory is unitary. To address this further, first and foremost, our model certainly exists by construction, and as we showed, it has a regime with massless flows between conformal fixed points, which are more conventional flows. Although the CFT’s are non-unitary, these massless flows still fit in the paradigm that all flows begin and end at conformal fixed points. One should mention that in recent studies of massless flows between non-unitary minimal CFT’s in 2D, such flows are still consistent with a c-theorem Ravanini ; MussardoLG ; Tanaka ; KlebanovMM ; Negro , where here $c=c_{{\rm eff}}$, where $c_{{\rm eff}}$ is the effective central charge that appears in the thermodynamic Bethe ansatz. Furthermore, there are examples of non-unitary minimal models that flow to unitary ones. Returning to our model, if one is forced to understand the flows for all regimes of the couplings $g_{1},g_{3}$, then a cyclic regime where $Q$ is positive is an inescapable conclusion. At small coupling, this is just above the SU(2) symmetric diagonals. Secondly, for the 2D models with beta functions that are identical to those considered here and reviewed in the Appendix, an exact S-matrix was even proposed LeClairSierra , further justifying that these models are well-defined.
@@ VI Conclusions
We have computed the beta functions for our model to 3 loops, and showed that they do not spoil the main features of the RG flows based on the 1-loop approximation found previously in ALDolls . For SU(2) broken to U(1) we re-summed an infinite number of contributions and conjectured a non-perturbative beta function. This model has a line of fixed points, which are new non-unitary CFT’s in 4 spacetime dimensions. There exists RG flows between non-trivial CFT fixed points, and we computed the relation between the anomalous dimensions in the UV and IR, and furthermore identified some special points where these anomalous dimensions are rational. There also exists a regime where the flows do not begin nor end at a fixed point but in fact are cyclic, and we computed the period $\lambda$ of the flow in terms of the RG invariant $Q$. We argued that this circumvents the paradigm that all QFT’s begin or end at a fixed point as proposed in Schwimmer ; Polchinski , simply because the model is non-unitary.
@@
The analysis of the higher order RG flows in this paper points to the existence of some new non-unitary CFT’s in 4 spacetime dimensions. We computed some anomalous dimensions of relevant perturbing operators in the UV and based on massless flows found some special rational exponents corresponding to flows that arrive in the IR via operators of integer dimension $\mathfrak{D}>4$. If these calculations are correct, they are still far from a complete understanding of these new CFT’s. For instance, what are the primary fields and their correlation functions? It is interesting to observe that rational exponents can exist in both 2 and 4 spacetime dimensions, whereas in 3 dimensions, non-trivial exponents obtained by means of the epsilon expansion are highly non-perturbative and irrational WilsonFisher . This is perhaps due to there being no obvious analog to the $J^{a}\widetilde{J}^{a}$ factorization of marginal operators in $3D$, which would require $J^{a}$ to have dimension $3/2$ such as $\phi^{3}$, such that the marginal interaction goes as $\phi^{6}$. In 2D the obvious generalization of this factorization is as a product of left-right moving chiral conserved currents, which have dimension 1.
@@
It would be desirable to have a more fundamental understanding of the $\cal K$}-insertions for our modified metric on the Hilbert space, since in this article it was just introduced in order to obtain the interesting OPE (23) since this structure leads to interesting beta-functions. Based on our results on massless flows above, by analogy with the classification of massless flows between non-unitary CFT;s in 2D Tanaka , we suggest that $\cal K$} can be interpreted as a topological defect, analogous to Verlinde lines in 2D CFT. In the 4D context, these should be codimension 2 defects, namely two dimensional surfaces. It is well accepted that such defects can modify the critical behavior in significant ways. The massless flows we found along the critical line $g_{1}=0$ should then be viewed as constrained by the U(1) fusion category. From this viewpoint, the line of fixed points we found are analogous to the radius of compactification of a free boson in 2D, which also has a duality. The flows in Tanaka are based on a study of the $SU(2)_{q}$ quantum group modular fusion category where $q$ is a root of unity, which is known to be a reduction of the U(1) category from various viewpoints. Very recent work of Witten is also suggestive of a such an understanding, where he argues that bras and kets can differ by an operator, which is analogous to our $\cal K$}-inner product where kets have an operator $\cal K$} insertion, whereas bras do not.
The most unexpected result we found is that the beta function for our 4D model is identical to current-current interactions in 2D. This was established to 3-loops, and we argued that this extends to all orders, and thus conjectured the non-perturbative beta functions in (53). It would be very interesting if this can be generalized through a procedure to “lift” 2D CFT’s to 4D more generally. A recent work along these lines is lifts .
@@
There is some evidence for an underlying ${\rm SL}(2,\mathbb{Z})$ structure in our model. Above we showed that the higher order beta functions have a $g\to 1/g$ duality, like an S-duality $\tau\to-1/\tau$. T-duality, $\tau\to\tau+1$ typically corresponds to a periodicity due to a $\vartheta$-angle, as in supersymmetric Yang-Mills theory. The cyclicity of the RG required us to identify $g_{3}=\pm\infty$, endowing the 2-parameter coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder, which is a form of periodicity. This by itself does not establish an ${\rm SL}(2,\mathbb{Z})$ symmetry, since we have not precisely identified the elliptic modulus $\tau$ as a function of the 2 couplings. The role of ${\rm SL}(2,\mathbb{Z})$ is perhaps clearer for the fully anisotropic case with the beta functions (36) since integration of these 1-loop beta functions already leads to elliptic functions LeClairSierra . Furthermore the 2D models, with identical beta functions to our models, were proposed to have the S-matrix of Zamolodchikov Zelip built out of elliptic functions Elliptic . This S-matrix is characterized by a $\mathbb{Z}_{4}$ symmetry which is a subgroup of ${\rm SL}(2,\mathbb{Z})$, generated by the matrix $A=-i\sigma^{2}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Furthermore, it is known that this S-matrix transforms covariantly under ${\rm SL}(2,\mathbb{Z})$, meaning the S-matrix transforms with phase factors or gauge transformations that preserve the structure of the Yang-Baxter equation Baxter ; Belavin . In Elliptic , for the SU(2) broken to U(1) model, which is the cyclic sine-Gordon model reviewed in the Appendix, the latter can be obtained from certain limits of Zamolodchikov’s elliptic S-matrix, and this is perhaps a way of understanding what is left of the ${\rm SL}(2,\mathbb{Z})$ structure for our two parameter model.
VII Acknowledgements
@@
We wish to thank Luis Alvarez-Gaumé, Denis Bernard, Igor Klebanov, Cheng-Yang Lee, Michael Peskin, Germán Sierra and Takahiro Tanaka for discussions.
VIII Appendix: Comparison with 2D current-current perturbations
@@
It is instructive to compare the above 3-loop results to exactly solvable models in 2 spacetime dimensions which have identical beta functions, since this can provide insights into higher orders in perturbation theory, and more importantly, it provides a guide as to what the physical implications could be.
@@
These flows only exist for $1<\Gamma_{\scriptscriptstyle{\rm UV}}<2$, where $\Gamma_{\scriptscriptstyle{\rm UV}}=1$ corresponds to the free-fermion point of the sine-Gordon model. This model is integrable and the relation (83) agrees exactly with the thermodynamic Bethe-ansatz analysis in Fendley2 , and this provides a non-trivial check of the validity of our beta functions to all orders.